OpenStudy (anonymous):

Stuck on calculus homework question... Find the derivative of the function using the definition of derivative: g(t)= 5/sqrt t PLEASE HELP!!!

OpenStudy (weasel123):

g(t) = 7/sqrt(t) = 7•t^(-1/2) g'(t) = 7•(-1/2)•t^(-3/2) [use the formula (x^n)' = n•x^(n-1)] g'(t) = (-7/2)•t^(-3/2)

OpenStudy (anonymous):

ooooh ! Okay, now I get it. Thank you!

OpenStudy (weasel123):

np